OpenStudy (anonymous):

Given the equation y − 4 = 3/4(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (anonymous):

do you know what is slope-intercept form?

OpenStudy (anonymous):

I think of understand it not to much.

OpenStudy (anonymous):

it is of the form : y=mx+c

OpenStudy (anonymous):

on observing the slope-point form we see that on one side there should be only y.

OpenStudy (anonymous):

\[\large y-4=\frac{3(x+8)}{4}\] \[\large y=\frac{3(x+8)}{4}+4\] \[\large y=\frac{3x+24+16}{4}\] \[\large y=\frac{3x+40}{4}\] \[\large y=\frac{3x}{4}+\frac{40}{4}\] \[\large y=\frac{3x}{4}+10\]

OpenStudy (anonymous):

@wolf1234 clear??